OpenStudy (anonymous):

sove -4 sqrt x 2 3=-1

OpenStudy (anonymous):

could you write that a bit neater? doesn't look right

OpenStudy (anonymous):

\[-4 \sqrt x+2+3=-1\]

OpenStudy (anonymous):

subtract 5 from both sides -4*sqrt(x) = -6 divide by - 4 sqrt(x) = 3/2 square both sides x = 9/4

OpenStudy (anonymous):

how about this one

OpenStudy (anonymous):

\[x^5/6 \over x ^2/5\]

OpenStudy (anonymous):

@robz8

OpenStudy (anonymous):

\[\frac{ 5x^{3} }{6 }\]

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

are you trying to solve for x?

OpenStudy (anonymous):

divide them

OpenStudy (anonymous):

then that is the most simplified answer

OpenStudy (anonymous):

@eSpeX

OpenStudy (espex):

I agree with the first problem and looking at the second problem, \(\frac{\frac{x^5}{6}}{\frac{x^2}{5}}\), you can rewrite it as \(\frac{x^5}{6} * \frac{5}{x^2}\) and simplify. Since there is no '=' you cannot solve for 'x'. @robz8 didn't teach and lead you to getting the answer, but it is accurate.

OpenStudy (anonymous):

:P

OpenStudy (espex):

When you multiply the numerators and denominators you get \(\frac{5x^5}{6x^2}\), and you can then cancel x's to get the simplification.

OpenStudy (anonymous):

simplify (x^1/4)^4/9